Tazicef 500 mg IM stat. Drug available: Tazicef 1g in a vial. The label states to add 3 ml of sterile water in which when reconstituted the concentration will be 1 g = 3.6 ml. How many mls will be given?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

        Formula

             Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

    Here

          Desired dose(D) = 500 mg.

          Dose in hand(H) = 1000 mg (1g)

              Quantity(Q) = 3.6 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                   = 500/1000 * 3.6

                                                   = 0.5 * 3.6

                                                   = 1.8 ml.